OpenStudy (anonymous):

Create a polynomial that fits the following conditions. Degree 4, 2 positive real zeros, 0 negative real zeros, 2 complex zeros

OpenStudy (asnaseer):

are you aware of the fact that if a polynomial has a root say, \(r_1\), then \((x-r_1)\) will be one of its factors?

OpenStudy (asnaseer):

e.g. if it has a root of say, 5, then (x-5) will be one of its factors.

OpenStudy (anonymous):

Yes.

OpenStudy (asnaseer):

good. secondly, are you also aware that complex root always appear as pairs - each one being a complex conjugate of the other?

OpenStudy (asnaseer):

which means if (a+ib) is on the roots, then so is (a-ib)

OpenStudy (asnaseer):

*is one of the roots

OpenStudy (anonymous):

Yes.

OpenStudy (asnaseer):

good. now finally you are told that the polynomial must be of degree 4 - so it has 4 roots. two of them must be positive real roots - lets call them \(r_1\) and \(r_2\) and two of them must be complex - lets call them \(z=a+ib\) and \(\bar z =a-ib\)

OpenStudy (asnaseer):

so we know this polynomial must be a product of its factors and can be written down as:\[f(x)=(x-r_1)(x-r_2)(x-z)(x-\bar z)\]now all you need to do is pick values for these variables.

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

Yes.

OpenStudy (asnaseer):

good - so you can answer the question from here now - correct?

OpenStudy (anonymous):

Not really. I knew all that stuff already, but I don't know how to answer the question.

OpenStudy (asnaseer):

ok, I showed that the polynomial in question has to be of the form:\[f(x)=(x-r_1)(x-r_2)(x-z)(x-\bar z)\]so, why don't you first pick any two positive numbers to represent \(r_1\) and \(r_2\).

OpenStudy (asnaseer):

just give me any two positive real numbers that come to mind

OpenStudy (anonymous):

1 and 2 \[(x-1)(x-2)\]

OpenStudy (asnaseer):

perfect - now pick a complex number of the form (a+ib) so you need to pick values for 'a' and 'b'

OpenStudy (asnaseer):

I am /assuming/ you know that \(i=\sqrt{-1}\)

OpenStudy (asnaseer):

are you stuck?

OpenStudy (anonymous):

4+2i, 4-2i

OpenStudy (asnaseer):

ok, so now we know your polynomial is of the form:\[f(x)=(x-1)(x-2)(x-(4+2i))(x-(4-2i))\]

OpenStudy (asnaseer):

if we multiply the last two factors out then, because they are complex conjugates of one another, we will end up with something that does NOT contain any complex numbers.

OpenStudy (asnaseer):

\[(x-(4+2i))(x-(4-2i))=x^2-8x+20\]

OpenStudy (asnaseer):

do you understand that last step?

OpenStudy (anonymous):

\[(x^2-3x+2)(x^2-8x+20)\]

OpenStudy (anonymous):

?

OpenStudy (asnaseer):

yes, and that can be expanded out to the final polynomial as follows:\[f(x)=(x-1)(x-2)(x^2-8x+20)=x^4-11x^3+46x^2-76x+40\]although I believe it should be ok to just leave it in the form of:\[f(x)=(x-1)(x-2)(x^2-8x+20)\]

OpenStudy (asnaseer):

NOTE: you can also multiply this final polynomial by any real number without affecting the roots. so, for example, this is also a solution:\[f(x)=10(x-1)(x-2)(x^2-8x+20)\]

OpenStudy (asnaseer):

I hope that all made sense? :)

OpenStudy (anonymous):

Yes, thanks!

OpenStudy (asnaseer):

yw :)